OpenStudy (anonymous):

integral from -2 to 2 of 1/(x^2-4) i spit into 2 integrals and do trig sub x=2sec@ @=theta. I get to -1/2 lim t->-2+ of [ln|csc@+cot@|]from t to 0. But when i turn my @ back into x's i get ln[x/(sqrt(x^2-4) + 2/(sqrt(x^2-4)) and now 0 is undefined since it produces a negative under the sqrt... what do i do?

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}\frac{ dx }{ x^2-4 }\]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

left side=\[\frac{ 1 }{ 2 }\lim_{t \rightarrow -2+} \int\limits_{t}^{0}[-\ln|\csc (\theta)+\cot(\theta)]\]

OpenStudy (anonymous):

how did you get that anti derivative?

OpenStudy (anonymous):

ah the integral should go away my bad

OpenStudy (anonymous):

it was integral of csc@ at some point

OpenStudy (anonymous):

but should be [-ln|csc@+cot@] evaluated from t to 0

OpenStudy (anonymous):

using partial fractions you get \[\frac{1}{4}\int \frac{1}{x-2}+\frac{1}{x+2}dx\]

OpenStudy (anonymous):

hm i used trig sub was that incorrect?

OpenStudy (anonymous):

i dont see any cotangent, but i suck at these

zepdrix (zepdrix):

Oh interesting :O Your steps look fine edcion. I wasn't sure what the problem was until I threw it into Wolfram. The integral does not converge over that interval.

OpenStudy (anonymous):

i get the log, and no chance for it to converge

OpenStudy (anonymous):

i am asked to evauluate however -_-

OpenStudy (anonymous):

you cannot

OpenStudy (kainui):

Good luck

OpenStudy (anonymous):

the integral (clearly improper as you wrote) does not exist, i.e. that limit does not exist

OpenStudy (anonymous):

my problem is when i turn my thetas back into x's i get this : for left side only: \[-\frac{ 1 }{ 2 }\lim_{t \rightarrow -2+}[\ln|\frac{ x }{ \sqrt(x^2-4) }+ \frac{ 2 }{ \sqrt(x^2-4) } t \to 0\]

OpenStudy (anonymous):

but can the limit be -inf or positive inf?

OpenStudy (anonymous):

that last thing should be t to 0 in which case i have negative sqrt in the denominators of both when evaluating at 0

OpenStudy (anonymous):

i am confused about the method to get your integral, but i am not saying it is wrong i got \[\frac{1}{4}(\log(x-2)-\log(x+2))\] but in either form it should be clear that the limit as x approaches -2 does not exist also as x approaches 2 the integral does not converge

OpenStudy (kainui):

Sub x=i2tan(theta) Idm try it lol

OpenStudy (anonymous):

ug i don't even know how to use imaginary numbers in this shiz yet -______-

OpenStudy (anonymous):

ill try partial fractions

OpenStudy (anonymous):

it is really snappy with partial fractions integral still does not converge though

OpenStudy (anonymous):

right but i need to know then if it is positive or negative infinity (what the prof wants)

OpenStudy (zarkon):

evaluating the antiderivative satellite gave you will tell you that it diverges to -infinity

OpenStudy (anonymous):

lol so will a picture!

OpenStudy (zarkon):

not really

OpenStudy (anonymous):

lol first i can't draw the picture, second still need to do it algebraically

OpenStudy (zarkon):

if you knew it diverged first then a picture will tell you that it must be -infinity

OpenStudy (anonymous):

im trying now with partial fractions

OpenStudy (anonymous):

yes that is what i meant it is pretty clearly negative on that interval

OpenStudy (zarkon):

yes it is

OpenStudy (anonymous):

you know how to do partial fractions the snap way?

OpenStudy (anonymous):

hm a/(x+2) + b/(x-2)

OpenStudy (anonymous):

yeah and to find \(a\) do this: since in \(\frac{a}{x+2}\) \(x\) cannot be \(-2\) put your finger over the factor \(x+2\) is \[\frac{1}{(x+2)(x-2)}\] and replace \(x\) by \(-2\) you get \[\frac{1}{\cancel{(x+2})(-2-2)}=-\frac{1}{4}\]

OpenStudy (anonymous):

for \(b\) do the same thing, only put your finger over \((x-2)\) and replace \(x\) by \(2\) you get \[\frac{1}{(2+2)\cancel{(x-2)}}=\frac{1}{4}\]

OpenStudy (anonymous):

the left side so far is convergent

OpenStudy (anonymous):

since the ln's have absolute value

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\lim_{x\to -2^+}\ln(x+2)\] does not exist

OpenStudy (anonymous):

integral of 1/x is ln|x|

OpenStudy (anonymous):

what about the absolute values?

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

i was thinking that ln|x+2| as x approaches -2+ is zero

OpenStudy (anonymous):

@satellite73 if your thing was correct. 1/4(ln|x-2|-ln|x+2|) then it would be positive inf right? since the - in front of the ln makes -inf turn into positive inf

OpenStudy (anonymous):

because as x-> -2+ ln|x+2| -> -inf

OpenStudy (anonymous):

or ln|approaches 0| -> -inf